OpenStudy (anonymous):

Solve for x , where x is a real number √(3x+24)-√(x+21)=1

OpenStudy (anonymous):

x = -1

OpenStudy (anonymous):

@mathdood explain your self

OpenStudy (anonymous):

@mathdood didnt you were a scrub

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

the answer is 4

OpenStudy (anonymous):

@mathdood www.youtube.com/watch?v=xpF5wOcy7os

OpenStudy (anonymous):

oh shoot my bad I was not thinking at all I ignored the sqrts haha

OpenStudy (anonymous):

(x+21)= -1 +2(sqrt(3x+24)) - (sqrt(3x+24))^2 (x-3x+22-24)/2= sqrt(3x+24) (-2x-2)^2= (3x+24) 4x^2-8x+4 = 3x+24 4x^2-11x-20 = 0 so yeah 4 having trouble factoring the last part